### v1

# Contrastive Learning for Task-Independent SpeechLLM-Pretraining

Abstract
Large language models (LLMs) excel in natural language processing but adapting these LLMs to speech processing tasks efficiently is not straightforward. Direct task-specific fine-tuning is limited by overfitting risks, data requirements, and computational costs. To address these challenges, we propose a scalable, two-stage training approach: (1) A task-independent speech pretraining stage using contrastive learning to align text and speech representations over all layers, followed by (2) a task-specific fine-tuning stage requiring minimal data. This approach outperforms traditional ASR pretraining and enables the model to surpass models specialized on speech translation and question answering while being trained on only 10% of the task-specific data.
Contrastive Learning for Task-Independent SpeechLLM-Pretraining
Maike Züfle  and  Jan Niehues Karlsruhe Institute of Technology, Germany {maike.zuefle, jan.niehues}@kit.edu
footnotetext: https://github.com/MaikeZuefle/contr-pretraining
1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in translation and language understanding tasks (OpenAI et al., 2024; Dubey et al., 2024). However, their reliance on written text limits their application to real-world scenarios, where communication often occurs through speech. Extending LLMs to spoken language could enable them to reason about and process spoken content effectively.
A straightforward approach is to use a cascaded pipeline (Shen et al., 2023) where speech is first converted to text using an automatic speech recognition (ASR) model, and the text is subsequently processed by an LLM. While simple, this approach has limitations: it cannot be trained end-to-end, ASR errors can propagate through the pipeline, and critical paralinguistic information, such as speaker features or noise, is lost during transcription.
To address these issues, recent works have proposed SpeechLLMs (Wang et al., 2024a; Tang et al., 2024; Nguyen et al., 2024; Chu et al., 2024), which connect speech encoders to LLMs via a trainable projector in an end-to-end manner. However, training them effectively remains a key challenge due to the scarcity of labelled data for downstream tasks (Xu et al., 2023; Ünlü Menevşe et al., 2024).
Pretraining the SpeechLLM before finetuning on task-specific data can help overcome this challenge and reduce the need for large downstream datasets. A straightforward strategy is to use ASR data, training on the ASR task itself (Tang et al., 2024; Wang et al., 2024b). However, this might lead to overfitting to this pretraining task.
Figure 1: An overview of our two-stage training process. First we perform task-agnostic pretraining, by leveraging ASR data to calculate a contrastive loss over a selected set of layers. Then we finetune on task-specific data. The speech encoder and the LLM remain frozen (indicated by the snowflake), we only train the projector.
Our work introduces contrastive learning as a task-independent pretraining strategy for SpeechLLMs using ASR data. Contrastive learning aligns the representations of paired speech and text inputs while encouraging representations of unrelated inputs to remain distinct. We demonstrate that applying a contrastive loss to the representations after every layer and pretraining on just 400 hours of parallel speech-text data serves as an effective, task-agnostic foundation for SpeechLLMs.
To evaluate our approach, we finetune the pretrained model on three downstream tasks: ASR, speech translation (ST), and speech question answering (SQA). We compare its performance against that of traditional ASR pretraining and next-word prediction (NWP) pretraining using mixed speech-text input. The contrastive strategy not only surpasses the other methods but also matches the performance of task-specific state-of-the-art (SOTA) models on nearly all metrics. We also analyse combinations of the pretraining methods to explore potential complementary interactions.
Scaling up contrastive pretraining to 1,400 hours yields additional performance gains, with our models surpassing specialized models and current SpeechLLMs. Even in simulated low-resource scenarios, where only 10% of downstream data is used for finetuning, our method consistently outperforms existing SpeechLLMs.
Our SpeechLLM architecture is simple yet flexible, consisting of a speech encoder, a projector, and an LLM. Only the projector is trained, making the model parameter-efficient and enabling maximum flexibility in adapting to the ever-evolving landscape of LLMs and speech encoders. Moreover, this approach preserves the LLM’s text-to-text capabilities, ensuring its functionality across both speech and text-based tasks. An overview of our approach is provided in Fig. 1.
The main contributions of our paper are:
  1. 1.
    We propose contrastive pretraining for SpeechLLMs, achieving state-of-the-art results in ST and SQA.
  2. 2.
    We demonstrate the effectiveness of our pretraining method in low-resource settings.
  3. 3.
    We provide a comprehensive analysis of contrastive learning for SpeechLLMs and its interaction with other pretraining methods.
  4. 4.
    We show that contrastive pretraining preserves a model’s ability to recognise paralinguistic speech features.
2 Related Works
SpeechLLMs.
SpeechLLM architectures typically consist of three main components: a speech encoder, a pretrained language model, and a projector linking the two. Training strategies differ based on which components are finetuned.
Chu et al. (2024) adopt a comprehensive strategy by finetuning all components. Similarly, other works explore finetuning both the projector and the encoder, alongside applying Low-Rank Adaptation (LoRA; Hu et al. 2022) to the LLM (Wu et al., 2023; Gong et al., 2024). In contrast, some methods freeze the LLM entirely, focusing on training only the encoder and projector (Wang et al., 2024b; Held et al., 2024). Several other approaches focus on finetuning the projector while introducing LoRA layers to the LLM, leaving the speech encoder and LLM frozen (Peng et al., 2024; Tang et al., 2024). An even more lightweight approach freezes both the encoder and the LLM, finetuning only the projection layers (Wang et al., 2024a).
In this work, we use the latter approach, using a frozen speech encoder and LLM while finetuning only the projection layers. This preserves the text-to-text abilities of the LLM, while also being parameter efficient.
Text and Speech Alignment.
While ‘alignment’ has various meanings (Hämmerl et al., 2024), in this work, we use it to refer to the process of creating meaningfully similar representations for text and speech. One approach to achieve this alignment is to directly finetune on the desired tasks without explicitly modelling alignment (Fang et al., 2024). An alternative is to use a downstream task to pretrain the SpeechLLM before finetuning on desired tasks, for example pretraining on ASR (Tang et al., 2024; Wang et al., 2024b) or ASR variants (Wang et al., 2024a).
Other works achieve alignment by encouraging paired speech and text embeddings closer together using an appropriate loss term. For example, Held et al. (2024) use an L2 loss on speech and text embeddings in addition to a KL-divergence loss on the output distributions, and Chuang et al. (2020) use a cosine distance loss between paired embeddings. Optimal transport has also been used to align paired embeddings in encoder-decoder speech translation systems (Le et al., 2023; Tsiamas et al., 2024).
Contrastive training is another widely used approach, bringing parallel text and speech representations closer while pushing non-parallel representations apart. It has been used in knowledge distillation for spoken language understanding (Zhu et al., 2022; Cappellazzo et al., 2024), as an auxiliary loss for encoder-decoder ST systems (Ye et al., 2022; Ouyang et al., 2023), and for music (Manco et al., 2022). It has also been applied to classification tasks, such as emotion recognition (Sachidananda et al., 2022; Zhou et al., 2024).
Other approaches mix text and speech in the input to help models learn both modalities. For example, Fang et al. (2022) align the output distributions of speech and mixed text-speech inputs, while Nguyen et al. (2024) and Peng et al. (2024) mix speech and text sequences for next-word prediction and question-answering tasks.
In this work, we leverage the previous success of contrastive learning on speech models and apply it to SpeechLLMs. We show how contrastive learning can be applied successfully in this new scenario and combined with other losses. In contrast to other works, we experiment with using optimal transport for the contrastive loss function and apply the loss to several hidden layers of the SpeechLLM.
3 Methods
We propose contrastive learning as a task-independent pretraining method to align speech and text representations, leveraging abundantly available ASR data. We compare with two ASR-based pretraining baselines: traditional ASR pretraining and next-word prediction with mixed text-speech spans, which we also introduce as a pretraining approach in this paper.
Notation.
Let $\mathbf{t}\in\mathbb{R}^{H\times N}$ represent a text sequence with $N$ tokens, where each token is embedded into a $H$-dimensional space. The speech sequence is embedded into $M$ vectors, with $\mathbf{s}\in\mathbb{R}^{H\times M}$. Note that usually $M>N$.
3.1 Contrastive Training
Contrastive training brings representations of matched speech-text pairs closer whilst pushing representations of mismatched pairs apart.
Harnessing both matched and mismatched pairs is often a very efficient way of using a dataset and allows one to leverage abundantly available ASR data.
In this paper, we employ the InfoNCE loss (van den Oord et al., 2019) similar to other works (Manco et al., 2022). InfoNCE loss is defined as:
$\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp\left(\frac{\text{sim}(\mathbf{s}_{i},\mathbf{t}_{i})}{\tau}\right)}{\sum_{j\in B}\exp\left(\frac{\text{sim}(\mathbf{s}_{i},\mathbf{t}_{j})}{\tau}\right)},$ (1)
where $\text{sim}(\cdot,\cdot)$ denotes a similarity measure, $\mathbf{s}_{i}$ and $\mathbf{t}_{i}$ denote speech and text embeddings respectively for the $i$-th example in the batch, and $\tau$ is the temperature parameter controlling the sharpness of the similarity scores. Summing over $B$ in the denominator includes the aligned speech-text pair $(\mathbf{s}_{i},\mathbf{t}_{i})$, as well as negative (contrastive) pairs. We choose to use all contrastive pairs within a mini-batch to allow for an efficient implementation.
The contrastive loss can also be calculated for multiple layers throughout the network and then summed. Fig. 1 illustrates this setup. We find that this gives the best results (see Section 5).
We adopt two different similarity measures:
Cosine Similarity.
One way to compute $\text{sim}(\mathbf{s}_{i},\mathbf{t}_{i})$ is to use cosine similarity. Since the length of the speech sequences, $M$, is typically much larger than the length of the text sequences, $N$, we cannot directly compute cosine similarity on the entire sequence. Instead, we average the embeddings over the sequence length to obtain a single embedding for each (Ye et al., 2022). We denote this pretraining strategy with contr-cos.
Wasserstein Distance.
To avoid aggregating the sequences, we can leverage the Wasserstein (or Optimal Transport) distance between two sequences (Peyré and Cuturi, 2019; Le et al., 2023). In our case, the Wasserstein distance involves finding the optimal way to "transport" mass from one normalized set of embeddings to another, enabling us to compute the alignment between text and speech embeddings of differing lengths.
More precisely, the optimal transport distance is calculated as follows: a mass of $1/N$ is placed at each text embedding location (points in a $H$-dimensional space) and we want to move this total mass of $1$ to the $M$ speech embedding locations, $1/M$ to each. There is a cost $C_{ij}$ associated with moving a unit of mass from location $t_{i}$ to location $s_{j}$ for all $i\in\{1,\ldots,N\}$ and $j\in\{1,\ldots,M\}$. The cost is given by the $L^{p}$ distance in the embedding space. This can be written as:
[…]
$\textbf{Z}\mathbf{1}^{N}=\mathbf{1}^{N}/N,\textbf{Z}^{T}\mathbf{1}^{M}=\mathbf{1}^{M}/M,\textbf{Z}\geq\textbf{0},$ (2)
[…]
Using the Wasserstein distance in the loss is expensive, so we resort to an upper-bound approximation that can be efficiently computed using the Sinkhorn algorithm (Sinkhorn and Knopp, 1967).
We use the negative distance as the similarity measure for $\mathcal{L}_{\text{InfoNCE}}$ and denote this pretraining strategy with contr-wasser.
3.2 Other Pretraining Approaches
We evaluate our contrastive pretraining by comparing it to traditional ASR pretraining and a mixed text-speech NWP pretraining approach.
ASR Pretraining.
We use task-specific fine-tuning on the ASR task, i.e. following the ASR-instruction prompt, for pretraining (Tang et al., 2024; Wang et al., 2024b) and use it as a baseline.
Mixed Next Word Prediction.
Inspired by Nguyen et al. (2024), who add mixed NWP to their finetuning data mix for downstream tasks,
we propose mixed text and speech NWP as a pretraining strategy.
We mix text and speech embeddings across randomly selected spans in the sequence, maintaining coherence with word boundaries. A sample sequence might look as follows: $t_{0}^{0},t_{1}^{0},s_{2}^{1},s_{3}^{1},s_{4}^{2},s_{5}^{2},s_{6}^{2},t_{7}^{3}$, where superscripts represent word boundaries, and subscripts represent subwords within the text or speech embedding. Since speech embeddings typically differ in length from text embeddings, we compute the NWP loss exclusively on text tokens. Details on how we select the speech and text spans are provided in Appendix A.
We also experiment with using these mixed sequences for contrastive learning: instead of aligning speech and text, we align mixed text-speech representations with text representations.
4 Experiments
Our experiments follow a two-stage training strategy: first, we perform text-speech alignment pretraining using ASR data to adapt the model to handle speech. In the second stage, we finetune the model jointly on a mix of task-specific data. This approach leverages abundant ASR data to reduce reliance on large task-specific datasets, enabling effective adaptation even in low-resource settings.
4.1 Model Architecture
Our SpeechLLM architecture consists of three core components: a speech encoder, a text-based LLM, and a projector bridging the two. We select HuBERT (Hsu et al., 2021) as our speech encoder due to its ability to process audio inputs longer than 30 seconds, which is crucial for SQA tasks. Recent research (Hassid et al., 2024; Nguyen et al., 2024) has demonstrated strong performance with HuBERT variants. We specifically utilize the facebook/hubert-large-ls960-ft variant. We use Llama-3.1-8B-Instruct (Dubey et al., 2024) as our LLM. As a projector, we choose Q-Former (Li et al., 2023), a window-level query Transformer, which already proved successful for other SpeechLLMs (Tang et al., 2024; Held et al., 2024).
For contrastive learning, we compare speech and text embeddings without the use of system prompts. This creates a positional mismatch between contrastive pretraining and finetuning. To address this potential overfitting problem, we experiment with varying the absolute starting position in the positional embeddings through random sampling for each example during pretraining.
4.2 Training Parameters
In both pretraining and task-specific finetuning, we keep the LLM and speech encoder frozen, adjusting only the speech projector, which has 42.5 million parameters. Training is conducted on four NVIDIA A100-SXM4-40GB GPUs. ASR pretraining takes approximately 9h, contrastive pretraining with cosine similarity over all layers takes around 5h. The task-specific finetuning takes around 30h with 100% of the data. Detailed training parameters and model specifics are listed in Appendix B.
4.3 Data
Pretraining.
For pretraining, we use the English portion of the widely-used MustC-v1 (Di Gangi et al., 2019) data (approximately 400 hours) to demonstrate effective performance with limited data. MustC includes TED talks with diverse speakers, accents, and topics, along with manually curated transcriptions. Additionally, we test pretraining on a larger dataset, adding the M-version of Gigaspeech (Chen et al., 2021) to our pretraining data, which includes 1,000 hours of speech from audiobooks, podcasts, and YouTube videos.
Finetuning.
We jointly finetune and evaluate our contrastive pretraining method on three diverse tasks covering a range of speech-text model abilities: Automatic Speech Recognition (ASR), Speech Translation (ST), and Speech Question Answering (SQA). For ASR and ST, we use the MustC-v1 dataset for training and testing. As before, English transcriptions serve as the ASR data, and we use language pairs en-de, en-fr, en-it, and en-es for ST. For question answering, we use Spoken-SQuAD (Lee et al., 2018), an English dataset based on Wikipedia, derived from the Stanford Question Answering Dataset (Rajpurkar et al., 2016, SQuAD).
Paralinguistic Feature Analysis.
We also analyse whether contrastive pretraining harms a model’s ability to capture paralinguistic features, such as speaking pace or noise. We use the mls-eng-speaker-descriptions dataset (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024), which involves reasoning about speaking rate, gender, background noise, and other features.
More information about the train and test sets can be found in Appendix C and details on prompts in Appendix D.
4.4 Evaluation
[…]
where $T$ is the set of tasks, and $score(t)$ is the model’s score on task $t$.
FT Model ASR ST SQA Contr loss on test Norm.
[…]
N/A Specialized 06.54 30.99 80.02 64.19 77.10 N/A N/A 100
N/A HuBERT + Llama 18.38 19.85 73.92 36.15 54.76 N/A N/A 0
100% no pretrain 12.09 28.84 79.94 62.39 74.44 1.37 1.26 79.96
10% no pretrain 23.78 19.87 69.72 31.35 41.86 1.38 1.36 -57.40
ASR pretrain 12.21 24.82 75.70 49.48 63.36 1.37 1.30 39.93
contr-cos-emb 13.98 25.47 76.32 55.10 68.72 0.97 0.72 46.34
contr-wasser-emb 15.99 26.10 77.09 55.04 68.17 1.08 0.62 44.02
contr-cos-all 13.06 27.19 78.29 60.48 73.90 1.08 0.91 67.39
contr-wasser-all 12.92 29.07 80.52 64.04 77.26 1.07 0.64 85.02
Table 1: Results of models with no pretraining, ASR pretraining, and contrastive pretraining using cosine similarity (contr-cos) or Wasserstein distance (contr-wasser) on embedding (-emb) or all (-all) layers, followed by fine-tuning on a 10% subset of task-specific data. The specialized baselines are Whisper (Radford et al., 2022) for ASR, Seamless (Communication et al., 2023) for ST, and You et al. (2022) for SQA. Metrics with * contribute to the normalized average as in Eq. 3.
(a) contr-cos pretrained models
(b) contr-wasser pretrained models
Figure 2: Analysis of a contrastive pretraining performed on different layers. We report the normalized average over tasks as in Eq. 3 for contrastive pretraining and finetuning on 10% of task-specific data. The last point on the x-axis refers to contrastive pretraining using the loss sum over all layers.
4.5 Baselines
We compare our approach against four baselines: (1) Firstly, we take three SOTA models, each specialized on one of the three tasks to set a goal for our models. Specifically, we use Whisper (Radford et al., 2022) for ASR, Seamless (Communication et al., 2023) for ST, and You et al. (2022) for SQA [^1] More recently, Manakul et al. (2024) report SQA results for Gemini-1.5-Pro, which underperforms You et al. (2022).. These SOTA results also serve as an upper bound for normalizing the respective scores when calculating the overall (normalized average) performance scores. (2) As the second baseline, we use the cascaded HuBERT-Llama-3.1-8B-Instruct model, which also serves as the lower bound for calculating overall performance scores.
Our primary objective is to assess the impact of contrastive pretraining. To do so, we need a controlled environment, where we keep the model architecture, data, and training setup as consistent as possible. To this end, we include two additional baselines using our training setup: (3) No pretraining and (4) Standard ASR pretraining.
5 Results
The goal of our experiments is to find the best setup for using contrastive learning for SpeechLLMs. We investigate various contrastive pretraining losses and where to apply them, combinations with other pretraining losses, and the impact of additional data. We then compare our best-performing model to recent SpeechLLMs. Finally, we investigate how text-speech alignment affects the detection of paralinguistic features.
5.1 Contrastive Pretraining
Contrastive Loss on the Embedding Layer.
We apply contrastive pretraining to the embedding layer using either Wasserstein loss (contr-wasser-emb) or cosine similarity loss (contr-cos-emb). Models are then jointly finetuned on three downstream tasks using 10% of the data to simulate low-resource settings (Table 1).
Our contrastive pretraining models consistently outperform the cascaded Hubert+Llama and no-pretraining baselines across all metrics. In ST and SQA tasks, they also surpass standard ASR pretraining, by over 5% in SQA. However, ASR-pretrained models remain better for ASR tasks. Interestingly, while contrastive models exhibit lower test-set contrastive loss on the embedding layer, the losses for no-pretraining and ASR-pretraining are similar despite a large performance gap.
Contrastive Loss on Different Layers.
Keeping with the low-resource setting, we investigate calculating the contrastive loss at different LLM layers. Figure 2 demonstrates that applying contrastive loss to deeper layers significantly increases downstream performance, with improvements of up to 20% for cosine similarity loss and 40% for Wasserstein loss.
Further analysis reveals that deeper layers benefit ST and SQA the most, while improvements for ASR are less consistent. We analyse both contrastive loss and task performance across layers in detail in Appendix E. To address positional mismatches between pretraining and finetuning (detailed in Section 4.1), we experiment with position shifts during pretraining. However, these do not consistently improve performance.
Contrastive Loss on All Layers.
Although Fig. 2 shows a clear advantage to applying the contrastive loss to deeper layers, the best layer to use is not consistent. Therefore, we also explore using a summed loss over all (multiples of 5) layers in Table 1. The methods contr-cos-all and contr-wasser-all achieve normalized task averages of $67.39$ and $85.02$, respectively, significantly outperforming the ASR-pretraining baseline and the contr-emb models. Contr-wasser-all, despite using only 10% of the finetuning data, even matches the performance of the specialized models on two out of three tasks and surpasses the no pretraining model finetuned on 100% of the task-specific data. This highlights the suitability of contrastive pretraining for scenarios with limited task-specific data.
5.2 Mixed-Text-Speech Input
Model ASR ST SQA Norm.
[…]
no pretrain 23.78 69.72 41.86 -57.40
ASR pretrain 12.21 75.70 63.36 039.93
contr-cos-emb 13.98 76.32 68.72 046.34
   + mixed 14.26 76.61 68.40 046.61
contr-wasser-emb 15.99 77.09 68.17 044.02
   + mixed 14.70 75.05 61.48 026.53
contr-cos-all 13.06 78.29 73.90 067.39
   + mixed 17.63 73.83 55.12 002.14
mixed-nwp 15.27 76.05 67.68 0 39.67
Table 2: Comparison of models pretrained on mixed speech-text input and finetuned on 10% of the task-specific data.
The previous section shows that contrastive learning is an effective pretraining strategy, outperforming ASR pretraining. We now analyse, whether aligning mixed speech-text and text sequences can further enhance performance.
Mixed Contrastive Pretraining.
We adapt our approach to use mixed speech-text sequences (contr-*-*+mixed). Using the contr-cos loss on the embedding layer, our model matches the performance of standard contrastive pretraining, outperforming both no pretraining and ASR pretraining (Table 2). However, with the contr-wasser loss results degrade significantly. Consequently, we continue with contr-cos-all+mixed to test the loss across all layers. This model also significantly underperforms its non-mixed counterpart. Further ablations of contrastive pretraining with mixed speech-text inputs do not yield additional improvements (detailed results, including BLEU and EM scores, are provided in Appendix F).
Mixed NWP.
We also propose using a next-word prediction (NWP) objective with mixed speech-text inputs (mixed-nwp) for pretraining, inspired by Nguyen et al. (2024). While this approach surpasses the no pretraining and ASR pretraining baselines (Table 2, last row), it remains inferior to contrastive pretraining.
5.3 Combining losses
Model ASR ST SQA Norm.
[…]
ASR pretrain 12.21 75.70 63.36 39.93
contr-cos-emb 13.98 76.32 68.72 46.34
   + asr loss 13.18 76.89 69.23 52.47
contr-wasser-emb 15.99 77.09 68.17 44.02
   + asr loss 13.19 77.4 70.79 57.54
contr-cos-all 13.06 78.29 73.90 67.39
   + asr loss 11.68 78.94 76.40 78.58
contr-wasser-all 12.92 80.52 77.26 85.02
   + asr loss 11.23 80.26 77.57 88.80
Table 3: Comparison of models pretrained on a combination of contrastive and ASR losses and then finetuned on a 10% subset of task-specific data.
FT Model ASR ST SQA Contr loss on test Norm.
[…]
N/A Specialized 06.54 30.99 80.02 64.1 77.1 N/A N/A 100
N/A BLSP-lslm-7b 44.51 28.70 78.68 5.60 21.82 N/A N/A -96.72
N/A Qwen2-Audio-7b 12.03 21.57 74.81 27.79 46.75 N/A N/A 010.79
10% ASR pretrain 12.21 24.82 75.70 49.48 63.36 1.37 1.30 039.93
   + giga 12.16 27.44 78.38 62.23 73.40 1.37 1.31 069.72
contr-cos-all 13.06 27.19 78.29 60.48 73.90 1.08 0.91 067.39
   + giga 10.94 29.95 81.21 65.58 79.32 1.06 0.71 097.43
contr-cos-all + asr 11.68 27.61 78.94 64.16 76.40 1.08 0.91 078.58
   + giga 11.12 29.90 81.29 72.43 82.63 0.98 0.65 102.30
100% ASR pretrain 10.28 30.31 80.98 70.06 81.01 1.37 1.22 100.53
   + giga 11.48 30.46 81.03 72.41 82.73 1.37 1.23 100.02
contr-cos-all 12.56 30.69 81.48 71.93 82.70 1.25 1.02 099.36
   + giga 9.31 31.36 81.98 75.28 84.74 1.10 0.83 114.29
contr-cos-all + asr 10.04 30.82 81.56 72.57 82.72 1.19 1.03 106.94
   + giga 10.03 31.54 81.99 76.25 85.01 1.10 0.79 112.76
Table 4: Comparison of pretraining only on Must-C data or the combination with the bigger Giga dataset (+giga). We compare our models against BLSP (Wang et al., 2024a) and the Qwen2-Audio-7B-Instruct model (Chu et al., 2024). Metrics with * contribute to the normalized average as in Eq. 3.
We now explore the potential of combining the previous pretraining objectives.
Contr. + ASR Loss.
We find that combining the ASR pretraining loss with the contrastive losses on the embedding layer leads to improved performance for ASR and SQA, as seen in Table 3. Similarly, this enhances the performance of the contr-*-all models, closing the gap between our contrastive pretrained models and ASR pretraining on ASR.
Contr. + mixed-NWP Loss.
In contrast, combining the mixed-NWP loss with the contrastive losses does not yield further improvements. Detailed results are provided in Appendix G.
5.4 Additional pretraining data
In the previous sections, we use Must-C ASR data for pretraining. Given the abundance of ASR data, we expand our pretraining dataset by incorporating GigaSpeech (Chen et al., 2021), effectively tripling the amount of data. The results demonstrate the significant impact of larger-scale pretraining.
Low-resource setting.
As before, we use the 10% finetuning subset to simulate the low-resource scenario. In preliminary experiments with lightweight contrastive pretraining on the embedding layer (Appendix H), we find that increasing the pretraining data leads to much larger improvements in the normalized average with contr-cos loss ($+33$) than with contr-wasser loss ($+6.10$). Therefore, we use contr-cos for subsequent experiments, where we perform contrastive learning on all layers with and without ASR loss, as these settings yielded the best results in the previous sections. As shown in Table 4, both models contr-cos-all+giga and contr-cos-all+asr+giga achieve an overall improvement of more than $20$ points compared to the same models trained with less pretraining data, reaching $97.43$ and $102.3$ points respectively. Remarkably, despite being finetuned on only 10% of the task-specific data, these models surpass both specialized models, the Seamless model (Communication et al., 2023) for ST and the You et al. (2022) model for SQA. Moreover, they outperform the ASR-pretraining baseline even on the ASR task.
Benchmarking our best variants.
We also evaluate our best-performing models in a higher resource setting. We finetune contr-cos-all+giga and contr-cos-all+asr+giga on 100% of the task-specific data. As expected, this leads to even stronger results. But while the ASR loss provides significant improvements in the low-resource settings, when finetuning with 100% of the available data, the contrastive loss alone performs on par with the combined approach. We hypothesize that the ASR loss is more sensitive to the out-of-domain Giga pretraining data than the contrastive loss, leading to no improvement on the ASR task.
We compare these models against blsp_lsm_7b (Wang et al., 2024a), which has a similar architecture to ours in that it trains only the projector while keeping other components frozen, and Qwen2-Audio-7b-Instruct (Chu et al., 2024), a strong baseline SpeechLLM. Notably, BLSP is finetuned on Must-C translation data, whereas for all other tasks, BLSP and Qwen2-Audio results reflect zero-shot performance. Across all tasks, our model outperforms both of these SpeechLLMs.
5.5 Impact of text-speech alignment on capturing paralinguistic features
Speech data is much richer than a sequence of words, and the pretrained speech encoder (HuBERT) might capture additional paralinguistic features. By aligning the richer speech embeddings more closely with text embeddings, the pretraining methods we explore could potentially train the projector to discard these additional features. To assess whether this occurs, we finetune our models on paralinguistic classification tasks before and after pretraining. In addition, we use the speech encoder with a linear classification head as a baseline.
The results demonstrate that models with contrastive and ASR-based pretraining perform comparably to the plain HuBERT model in predicting paralinguistic annotations, indicating that alignment-focused pretraining does not significantly affect the retention of paralinguistic information. Detailed accuracy results are presented in Table 12 in Appendix I. In addition, Table 13 shows that incorporating these features as auxiliary finetuning tasks does not notably alter model performance.
6 Conclusion
We introduce a contrastive learning-based pretraining strategy for SpeechLLMs, aligning text and speech embeddings in a parameter-efficient manner. This approach establishes an effective task-agnostic foundation for SpeechLLMs and demonstrates state-of-the-art performance across multiple downstream tasks, even when little finetuning data is available.
Importantly, this pretraining method preserves the model’s ability to understand paralinguistic information, offering a promising path forward for advancing SpeechLLMs.
7 Limitations
While our work provides valuable insights into pretraining of LLMs, it has some limitations that offer opportunities for future research:
More Expensive Training Scenarios:
Our study focuses on a lightweight and flexible approach, keeping the LLM and speech encoder frozen while fine-tuning only the projector. Although this minimizes computational costs, future research could explore scenarios where the LLM and encoder or adapter layers are also fine-tuned, potentially unlocking additional capabilities.
Limited Incorporation of Meta-Speaker data:
We analyze whether contrastively pretrained models, designed to align speech and text, retain their ability to capture meta-speech information such as speaker characteristics or noise. While we find that these models do not lose this capacity, extending this analysis to generation tasks where such meta-speech features influence the output would provide a deeper understanding of their impact.
Low-resource Scenarios:
Our study simulates low-resource settings by using subsets of the fine-tuning data, finding that models perform well even with only 10% of the task-specific data. However, extending this analysis to truly low-resource languages or tasks would offer richer insights into the model’s adaptability and performance under resource constraints.
Analysis of Contrastive Examples:
In this work, we use all contrastive pairs within a minibatch to maintain an efficient implementation. However, an alternative approach could involve selecting contrastive examples that are more challenging by ensuring they are closer to the positive examples, as demonstrated by Ye et al. (2022).
Potential Risks
Enabling Language Models to understand speech makes them more easily and widely accessible. However, this also lowers the barrier for the potential misuse of these models. In addition, SpeechLLMs can reproduce biases seen in the training data. We try to minimize this risk by not training the LLM, so any safeguard measures trained into the LLM also apply to the SpeechLLM. However, the model could still be affected by adversarial attacks.
Acknowledgments
This work has received funding from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People). We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid ACC Cyfronet AGH for providing computer facilities.
Appendix A Mixed Next Word Prediction
For alignment between text and speech, we utilize the forced alignment algorithm described by Kurzinger et al. (2020), which leverages wav2vec 2.0 (Baevski et al., 2020).
We then randomly select spans from a sequence and assign them either text or speech embeddings, ensuring that sequences begin with text or speech equally often. Because speech embeddings are generally longer than text embeddings for the same content, we select spans of 2 to 5 words for speech and 4 to 10 words for text, inspired by Nguyen et al. (2024) who use a similar distribution. However, unlike Nguyen et al. (2024), we do not introduce modality-specific tokens, as their framework uses these tokens to predict speech output — a capability we do not require.
In the alignment process for speech and text, the aligner generates timestamps for each word. If there is a pause between words, this pause is typically excluded from the alignment. However, to prevent the model from encountering unexpected silence or noise before a word during fine-tuning, we include any preceding silence or noise in the audio segment before the word.
Appendix B Hyperparameters and Model Details
All hyperparameters, model and training parameters are listed in Table 5.
Our SpeechLLM architecture consists of three core components: a speech encoder, a text-based LLM, and a projector bridging the two. We select facebook/hubert-large-ls960-ft (Hsu et al., 2021) as our speech encoder, Llama-3.1-8B-Instruct (Dubey et al., 2024) as our LLM and Q-Former (Li et al., 2023) as the projector.
HuBERT has an apache-2.0 license, the llama model has a llama3.1 license. We adhere to both license requirements, when using the models for research.
For contrastive learning, we compare speech and text embeddings without the use of system prompts. This creates a positional mismatch between contrastive pretraining and fine-tuning. When using ROPE embeddings (Su et al., 2023), this impacts contrastive pretraining at layers beyond the embedding layer. To address this potential overfitting problem, we experiment with varying the absolute starting position in RoPE by randomly sampling it during pretraining for each example.
training Q-Former Num Query Token 4
parameters Q-Former Num Hidden Layers 4
Q-Former Num Attention Heads 13
Q-Former Seconds per Window 1/3
per device batch size 10
gradient accumulation steps 2
num GPUs 4
learning rate 1e-4
warmup ratio 0.03
optimizer adamw_torch
learning rate scheduler type cosine
model max length 2048
gradient clipping 1
pretraining num epochs 5
specific contrastive $\tau$ cos + wasser 0.1
contrastive $\tau$ nwp 0.5
sinkhorn loss p 2
sinkhorn loss blur 0.5
finetuning num epochs 2
Table 5: Hyperparameters for the trainings, which are conducted on four NVIDIA A100-SXM4-40GB GPUs, mostly following Verdini et al. (2024)
Appendix C Data
We fine-tune and evaluate our contrastive pretraining method on three task: Automatic Speech Recognition (ASR), Speech Translation (ST), and Speech Question Answering (SQA).
For ASR and ST, we use the MustC v1 dataset (Di Gangi et al., 2019). We use the en-de portion as English ASR data, and we use language pairs en-de, en-fr, en-it, and en-es for ST. We exclude examples longer than 45 seconds for training and use the tst-common and tst-he test sets for evaluation.
For question answering, we use Spoken-SQuAD (Lee et al., 2018), which is based on Stanford Question Answering Dataset (Rajpurkar et al., 2016, SQuAD). To allow for higher batch sizes during training, we exclude examples exceeding 150 seconds.
For pretraining, we use the MustC ASR data (approximately 400 hours), but also experiment with adding the M-version of Gigaspeech (Chen et al., 2021), which includes 1,000 hours of speech from audiobooks, podcasts, and YouTube.
Contrastive pretraining aligns text and speech, which could potentially lead to the loss of paralinguistic features in the embeddings. Therefore, we also train and test our models on mls-eng-speaker-descriptions (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024). We use the mls-eng-speaker-descriptions train and test set. The available training data is quite large, so we subsample it to make training feasible. Since the test set is predefined, we sample the training data such that the test data size is 5% of the data. We let the model predict the following features: Speaking rate, gender, noise, reverberation, speech monotony, SDR noise, and PESQ speech quality.
Details on prompts can be found in Appendix D.
Appendix D Prompts
In this sections we report the prompt we used for the different tasks during inference. During training, for each example, the prompt is sampled from a list of multiple prompts. Here, we report the prompts used for testing, for the training prompts, we refer the reader to our codebase.
Task Prompt
ASR Can you transcribe this audio?
SQA Listen to the audio and answer this question:
Table 6: ASR and SQA prompts used for inference. We finetuned on a range of different prompts, please refer to our codebase for these.
Prompt
en Can you translate this audio from <sl> into <tl>?
fr Pouvez-vous traduire cet audio de <sl> en <tl>?
it Puoi tradurre questo audio da <sl> in <tl>?
es ¿Puedes traducir este audio de <sl> al <tl>?
de Können Sie dieses Audio von <sl> zu <tl> übersetzen?
Table 7: ST prompts used for inference. We finetuned on a range of different prompts, please refer to our codebase for these. For <sl> the source language is added, for <tl> the target language is added.
Prompt
Evaluate the audio file and provide details in the following order: speaking rate, gender, noise level, reverberation, speech monotony, SDR noise, and PESQ speech quality. Format the response as: <speaking_rate>, <gender>, <noise>, <reverberation>, < speech_monotony>, <sdr_noise>, <pesq_speech_quality>.
Table 8: Prompts for the MLS speaker description dataset used for inference. We finetuned on a range of different prompts, please refer to our codebase for these.
(a) Loss during contr-cos-petraining
(b) Loss during contr-wasser-petraining
Figure 3: Contrastive loss during pretraining performed on multiple layers for models trained with contrastive average loss (left) and contrastive wasserstein loss (right).
(a) ASR: contr-cos pretrained
(b) ASR: contr-wasser pretained
(c) ST: contr-cos pretrained
(d) ST: contr-wasser pretained
(e) SQA: contr-cos pretrained
(f) SQA: contr-wasser pretained
Figure 4: Analysis of a contrastive pretraining performed on multiple layers for different tasks. The performance is reported after finetuning on 10% of the task specific data. The last point on the x-axis refers to contrastive pretraining using the loss sum over all layers.
Appendix E Contrastive Loss over Multiple Layers
This section analyses contrastive pretraining over multiple layers. Fig. 3 shows the contrastive loss in the pretraining phase. The later the layer, the lower the contrastive loss, with an exception of the embedding layer. For cosine contrastive training, these differences are significanlty bigger than for wasserstein contrastive training. Fig. 4 shows the finetuning performance on 10% on task-specific data after pretraining on different layers for the tasks ASR, ST and SQA. Here the losses for pretraining seem to correlated mostly with the performance of the model after finetuning: Models perform better after contrastive pretraining at later layers. However, for ASR, this trend is not so noticeable.
FT Model ASR ST SQA Norm.
Data WER* BLEU COMET* EM F1* avg.
10% no pretrain 23.78 19.869 69.722 31.345 41.859 -57.401
ASR pretrain 12.21 24.822 75.702 49.478 63.358 39.934
10% contr-cos-emb + mixed 14.26 25.96 76.605 51.968 68.395 46.614
contr-wasser-emb + mixed 14.7 24.145 75.045 48.012 61.48 26.531
mixed-contr-cos-last 19.02 23.023 73.92 46.285 58.557 3.859
mixed-contr-wasser-last 16.03 23.09 74.14 43.835 58.201 12.947
contr-cos-all + mixed 17.63 22.901 73.828 42.149 55.117 2.136
10% mixed-nwp with punctuation 15.76 24.229 75.085 47.651 61.36 23.586
mixed-nwp without punctuation 15.27 25.224 76.05 53.574 67.678 39.667
mixed nwp without punctuation + giga 14.91 26.071 77.172 56.988 70.935 51.673
Table 9: Ablation studies for models with mixed speech and text input. Metrics with * contribute to the normalized average as in Eq. 3.
Appendix F Mixed Speech-Text Inputs
We experiment with mixed speech-text input as described in Section 3.2, and report results in Table 9.
We test contrastive pretraining with mixed speech-text input (contr-*+mixed) and find that aligning the representation on the embedding layer works best (however, contrastive learning without mixed input still works better). We also experiment with mixed speech-text next word prediction (nwp). Here, removing punctuation seems to help the model understand the NWP task better. Finally, we also combine the NWP task with calculating the contrastive loss and summthe two losses, which also does not seem to help the model to align text and speech embeddings.
FT Model ASR ST SQA Norm.
[…]
10% ASR pretrain 12.21 24.822 75.704 49.478 63.358 39.934
10% contr-cos-emb 13.98 25.474 76.322 55.1 68.719 46.339
   + asr loss 13.18 25.773 76.892 56.124 69.229 52.467
   + mixed + nwp 14.68 25.282 76.175 48.855 63.922 36.406
contr-wasser-emb 15.99 26.096 77.085 55.04 68.166 44.024
   + asr loss 13.19 26.552 77.4 59.016 70.79 57.544
   + mixed + nwp 16 23.98 74.593 45 59.113 16.868
10% contr-cos-all 13.06 27.188 78.285 60.482 73.9 67.388
   + asr loss 11.68 27.608 78.94 64.157 76.396 78.577
   + mixed + nwp 22.08 23.16 74.195 48.474 62.126 2.073
contr-wasser-all 12.92 29.072 80.522 64.036 77.258 85.017
   + asr loss 17.63 29.064 80.26 66.727 77.565 70.784
   + mixed + nwp 15.72 24.13 74.993 51.205 64.254 27.515
Table 10: Comparison of models with combinations of alignment losses. Metrics with * contribute to the normalized average as in Eq. 3.
Appendix G Combined Losses
This section explores the combination of different losses. We combine 1) ASR and contrastive loss and 2) NWP and contrastive loss. The result can be found in Table 10. While 1) leads to signifiant improvements, 2) does not seem to work together.
Appendix H Adding more pretraining data
We analyize the effect of contrastive pretraining with more data, adding the Giga dataset to our pretraining data. Results for finetuning on 10% of the data can be found in Table 11.
FT Model ASR ST SQA Contr loss on test Norm.
[…]
N/A Specialized 06.54 30.99 80.02 64.1 77.1 N/A N/A 100
N/A BLSP-lslm-7b 44.51 28.70 78.68 5.60 21.82 N/A N/A -96.72
N/A Qwen2-Audio-7b 12.03 21.57 74.81 27.79 46.75 N/A N/A 010.79
10% ASR pretrain 12.21 24.82 75.70 49.48 63.36 1.37 1.30 039.93
   + giga 12.16 27.44 78.38 62.23 73.40 1.37 1.31 069.72
contr-cos-emb 13.98 25.47 76.32 55.10 68.72 0.97 0.72 046.34
   + giga 14.24 28.76 80.18 63.94 76.60 0.97 0.47 078.45
contr-wasser-emb 15.99 26.10 77.09 55.04 68.17 1.08 0.62 044.02
   + giga 16.36 26.47 77.59 58.05 71.09 1.06 0.58 050.12
contr-cos-all 13.06 27.19 78.29 60.48 73.90 1.08 0.91 067.39
   + giga 10.94 29.95 81.21 65.58 79.32 1.06 0.71 097.43
contr-cos-all + asr 11.68 27.61 78.94 64.16 76.40 1.08 0.91 078.58
   + giga 11.12 29.90 81.29 72.43 82.63 0.98 0.65 102.30
Table 11: Comparison of pretraining only on Must-C data or the combination with the bigger Giga dataset (+giga). We compare our models against BLSP (Wang et al., 2024a) and the Qwen2-Audio-7B-Instruct model (Chu et al., 2024). Metrics with * contribute to the normalized average as in Eq. 3.
Appendix I Impact of text-speech alignment of capturing paralinguistic features.
Speech data is much richer than a sequence of words, and the pretrained speech encoder (HuBERT) might capture additional paralinguistic features. However, contrastive and ASR pretraining could potentially lead to discarding these paralinguistic features, when pulling speech embeddings closer to text embeddings. To assess this, we finetune our models on paralinguistic classification tasks before and after pretraining. We also compare our models against a HuBERT + linear head baseline, to assess which features the encoder captures in the first place.
To this end, we use the mls-eng-speaker-descriptions train and test set (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024). This dataset contains classification tasks for rate, gender, background noise, and other features (more details in Appendix C).
The results demonstrate that models with contrastive and ASR-based pretraining perform comparably to the plain HuBERT model in predicting paralinguistic annotations, indicating that alignment-focused pretraining does not significantly affect the retention of paralinguistic information. Table 12 shows the accuracies for the different classification tasks for the different models.
In a second step, we asses whether adding paralinguistic features during training (again using the mls-dataset), influences the models performance on downstream tasks. We finetune on a 10% subset of finetuning data to simulate a low-resource scenario. We find that including the paralinguistic data does not lead to significantly different results, which is not surprising given that the tasks that we test do not rely on paralinguistic information. Results can be found in Table 13.
Model Speaking Gender Noise Reverberation Speech SDR PESQ
Rate Monotony Noise Speech Quality Avg.
random 14.29 50.00 14.29 20.00 20.00 16.67 16.67 22.25
HuBERT + Class. Head 57.38 64.69 27.89 28.12 39.35 42.85 27.22 41.07
no pretrain 57.07 64.26 27.62 27.38 37.60 33.86 23.72 38.79
ASR pretrain 57.23 57.95 28.87 29.05 36.35 30.19 21.86 37.36
contr-cos-all 56.46 62.17 28.60 29.27 35.39 34.17 24.12 38.60
contr-wasser-all 56.12 62.30 28.26 27.57 37.70 34.25 24.30 38.64
contr-cos-all + asr 55.00 61.29 28.73 26.24 35.29 30.94 24.86 37.48
contr-wasser-all + asr 56.86 59.62 28.31 29.21 36.64 40.25 26.37 39.61
Table 12: Accuracy in % results for different meta-speech categories on the mls-eng-speaker-descriptions dataset. Bold results are the best or not significantly worse than the best, underlined results are the weakest or not significantly worse than the weakest result using pairwise t-test ($p<0.01$).
FT Model ASR ST SQA Norm.
[…]
10% ASR pretrain 12.21 24.822 75.704 49.478 63.358 39.934
   + meta 12.11 24.91 75.702 49.472 63.31 40.154
10% contr-cos-all 13.06 27.188 78.285 60.482 73.9 67.388
   + meta 13.53 27.099 78.213 61.024 74.745 66.932
contr-wasser-all 12.92 29.072 80.522 64.036 77.258 85.017
   + meta 11.97 29.283 80.475 66.124 78.25 88.915
10% contr-cos-all + asr 11.68 27.608 78.94 64.157 76.396 78.577
   + meta 12.01 27.817 78.917 63.112 75.981 76.903
contr-wasser-all + asr 11.23 29.064 80.26 66.727 77.565 88.802
   + meta 11.16 29.342 80.547 65.402 78.26 91.604
Table 13: Comparison of models with and without pretraining, and with and without meta-speaker data in the finetuning process (on 10%). Metrics with * contribute to the normalized average as in Eq. 3.